OpenStudy (anonymous):

What is the approximate weight of a 20-kg cannonball on earth? A. 2 N B. 20 N C. 200 N D. 2000 N

OpenStudy (anonymous):

medal....

OpenStudy (anonymous):

weight = force = mass * acceleration

OpenStudy (anonymous):

mass* gravity =20*9.81=196.2 N ..... but to make the problem easy gravity=10.. so 10*20=200 N

OpenStudy (anonymous):

thxs:)